OpenStudy (anonymous):

Write the equation in standard form: y+3= 1/4 (x-1)

OpenStudy (shamil98):

y + 3 = (x-1)/4 multiply both sides by 4 4y + 12 = x -1 change to ax + by = c form -x + 4y = -11 or x - 4y = 11

OpenStudy (anonymous):

thank you